Hi sir,,please post the solution of the problem,


Q:The equation of the plane bisecting the acute angle between the planes 2x-y+2z+3=0 and 3x-2y+6z+8=0 is.........


Ans. 23x-13y+32z+45=0

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Tapasranjan

for acute angel bisector

(2x-y+2z+3)√(22+12+22)= - ( 3x-2y+6z+8)/√(32+22+62)

7(2x-y+2z+3) =-3(3x-2y+6z+8)

23x-13y+32z+45=0

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin